there are three numbers such that the second is three times the first nd the third is 3 less than four timed the first if the sum of the numbers 37 find
these numbers

Answer by RAY100(1637)

let x = 1st number
second number is (3*x)
third number is (4*x -3)
if sum is 37 then
37 = (x) + (3*x) + (4*x-3) = 8*x -3
add 3 to both sides
40 = 8*x
divide both sides by 8
x = 5

check ist (5) 2nd( 3*5) or (15), 3rd ( 4*5-3) or (17) , sum is 37 ok